write an equation that describes the relationship of the 2 variables. use your equation to finish the tables
x y
2 3
3 5
5 9
10 19
100
200
1

Answer by venugopalramana(3286)

x y
2 3
3 5
5 9
10 19...up to this it is a straight line..since we find
(3-2)/(5-3)=(5-3)/(9-5)=(10-5)/(19-9)=1/2
its eqn. is
y-3=2(x-2)




--------------------

100....what is this???
200....what is this???..
1.......what is this???check..